What is a subshell in quantum mechanics. my teacher is saying it has the same n value and l value. Is that true?
Answer #1

Not true

Subshell only depend upon of l .

l = 0 ( S Subshell )

l = 1 ( p Subshell )

l = 2 ( d Subshell )

Value of l = 0 to n-1 ( so n and l can never be equal )

Only naming of Subshell depend upon n , l but both value can never be same . Example

2p Subshell ---- n= 2 , l = 1